OpenStudy (anonymous):

5, 30, 155, 780, ... Please give the next three values in the sequence.

OpenStudy (anonymous):

look at the differences and you should see the pattern

OpenStudy (anonymous):

30 - 5 = 155 - 30 = 780 - 155 =

OpenStudy (anonymous):

HI :DDD